OpenStudy (anonymous):

what is the piont to 2x+3y=-5 5x-2y=-2 SHOW WORK :)

OpenStudy (anonymous):

you write it as a matrix and do row reduction. However, I assume you dont know how to do this, so ill do it a slightly different way which comes down to the same thing: the point is that you want to get an x and y so both equations make sense right? so here we go

OpenStudy (anonymous):

a way you can see it is that you define both equations a certain way and then and equate them to each other: so 2x + 3y =-5 can be written as 2x=-5-3y and 5x - 2y = -2 can be written as 5x= 2y-2 so if we multiply the first equation by 5 and the second by 2 we get 10x = -25 - 15y and 10x = 4y-4 as both equations are equal to 10x we can equate them to each other so -25-15y = 4y- 4 which can be written as 19y = -21 so y = -19/21. Now we know y we can plug it back into your equation

OpenStudy (anonymous):

h my bad y=-21/19 then work out x by plugging it back into one of your equations and check